OpenStudy (anonymous):

1) Russell deposited $650 in a savings account that earns 3% simple annual interest. If he makes no other deposits or withdraws, what will the balance of Russell's account be after 6 years? A) $117 B) $659 C) $767 D) $1,170

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

@kohai

OpenStudy (kohai):

So, you're going to use this formula: PV * (1+r)^n = FV Where PV is the present value ($650), r is the rate (3%), n is the time (6) So it's going to be set up like this: 650 * (1+.03)^5 = FV 650 * 1.03^5 = FV Then just continue and solve it to get the value